OpenStudy (anonymous):

HELP ASAP PLEASE!!! I learned percents a long time ago and I forgot how to do them. I want to go to a store and buy something that costs $19.99 and I have a 40% off coupon. How do I find 40% of $19.99?

OpenStudy (dan815):

40% means 40/100

OpenStudy (dan815):

of means multiplication in math

ganeshie8 (ganeshie8):

multiply 19.99 it wid 40 %

OpenStudy (dan815):

so 40% of 19.99 or 40% * 19.99 or 40/100 * 19.99

ganeshie8 (ganeshie8):

% = 1/100

OpenStudy (dan815):

like 3 of apples is llike saying 3* apples or 5 of bananas saying 5*bananas 40% of 19.99 is saying 40*19.99 = 40/100*19.99

OpenStudy (anonymous):

Ok...on my calculator I did 19.99*40%, I got 7.996. So $7.99 is 40% of 19.99?

OpenStudy (dan815):

yes

OpenStudy (dan815):